OpenStudy (anonymous):

solve each system by substitution. -3x-2y=-4 y=-7

OpenStudy (anonymous):

just put -7 in for y on the first equation. I'm sure you can do it.

OpenStudy (anonymous):

x=-3.33

OpenStudy (anonymous):

Show me what you did

OpenStudy (anonymous):

oaky

OpenStudy (anonymous):

-3x-2(-7)=-4 -3x+14=-4 -14 -14 -3x = 10 /-3 /-3 x= -3.33

OpenStudy (anonymous):

ok, -4-14 is the same as -4+-14.

OpenStudy (anonymous):

don forget keep-change-change. Do you know what that rule is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, here